Find the caterer's fixed cost and the cost per student served if the cost for 100 students will be 750 and the cost for 150 students will be 1050
Answer by richwmiller(17219)

Think of nunber of students as x and the total cost=y
fixed cost =b and cost per student =m
y=mx+b
So to write the equation we first find m the slope (cost per student)
and then find b the fixed cost
(y-y)/(x-x)=m
1050-750/150-100=m
300/50=6=m=cost per student
y=mx+b
1050=6*150+b
150=b
y=6x+150
150 is the fixed cost